# Theme-Explanation Structure for Table Summarization using Large Language Models: A Case Study on Korean Tabular Data

### v2

@@ Abstract
This paper proposes the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline designed to process tabular data efficiently. Tabular-TX preprocesses tabular data by focusing on highlighted cells. It then generates summary sentences following a structured format, where the Theme Part appears as an adverbial phrase, and the Explanation Part follows as a predictive clause. This approach enables tailored analysis by considering the structural characteristics of tables and their comparability. Unlike conventional fine-tuning approaches that require extensive labeled data and computational resources, our method leverages In-Context Learning to dynamically adapt to different table structures without additional training, ensuring efficient and scalable table interpretation. Experimental results demonstrate that Tabular-TX significantly outperforms conventional fine-tuning-based methods, particularly in low-resource scenarios, by leveraging table structures and metadata more effectively through structured prompts. The results confirm that Tabular-TX enables more effective processing of complex tabular data. Furthermore, it serves as a viable alternative for table-based question answering and summarization tasks in resource-constrained environments.
@@ 1 Introduction
footnotetext: * These authors contributed equally.footnotetext: $\dagger$ Corresponding author.
Tables are an essential visualization tool for concisely organizing and comparing complex data. However, tabular data is inherently compact, requiring higher analytical capabilities compared to text data. Current large language models (LLMs) face challenges in effectively handling the structural complexity of tabular data Liu et al. (2024). Therefore, enabling LLMs to process table data efficiently and perform Table Question-Answering (TableQA) remains a significant challenge.
Previous studies on TableQA primarily aimed to enhance LLM performance through fine-tuning methods Zhang et al. (2024b). However, fine-tuning requires substantial computational resources Strubell et al. (2019), and Korean tabular datasets are limited in size, making it challenging to fine-tune LLMs effectively. For example, studies leveraging table-related datasets such as FeTaQA Nan et al. (2022) have shown performance improvements through fine-tuning but suffer from high computational costs. Meanwhile, synthetic dataset generation techniques Hou et al. (2024) have been proposed to address data scarcity. However, repetitive training on generated data has been shown to degrade model performance Alemohammad et al. (2024).
This study aims to enhance LLMs’ in-context learning (ICL) ability to generate interpretations of table segments using the table interpretation corpus provided by the National Institute of Korean Language NIKL (2024). As shown in Figure 2, the NIKL task is to summarize the highlighted cells, which are labeled in the metadata as highlighted_cells, into a single coherent sentence. To achieve this, we propose the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline.
Tabular-TX simplifies the complexity of tabular data and conducts tailored analysis considering its structural characteristics. Specifically, it segments table summaries into two distinct components: the Theme Part, expressed as an adverbial phrase, and the Explanation Part, presented as a predicative clause. This structured approach ensures a more comprehensive analysis by considering the metadata and the cell data. Furthermore, we introduce a journalist persona into the LLMs to enhance the objectivity and clarity of the generated sentence.
@@
Our main contributions are as follows:
  • We propose the Tabular-TX, a method for generating table summaries based on the theme-explanation structure, significantly improving the interpretative performance of table segments.
  • We suggest preprocessing techniques to address complex structural issues in tabular data to improve LLM’s ability to interpret table data.
  • Our experimental results of Tabular-TX show high performance without fine-tuning by utilizing LLMs’ in-context learning (ICL) ability.
@@ 2 Related Work
2.1 Table Question Answering (TableQA)
@@
TableLlama Zhang et al. (2024b) aims to generalize table-based models beyond task-specific constraints by fine-tuning 14 datasets across 11 tasks, including Highlighted Cells QA. The model achieved performance comparable to or surpassing task-specific models and even outperformed GPT-4 OpenAI (2023) on unseen tasks. However, despite its effectiveness, the model suffers from high computational costs. The training process demands 48 NVIDIA A100 80GB GPUs, making fine-tuning infeasible in resource-constrained environments.
Among table-based interpretation benchmarks, FeTaQA Nan et al. (2022) serves as a key reference dataset for evaluating table summarization models. While Chain-of-Table and TableLlama utilize ICL Brown et al. (2020) and fine-tuning, respectively, they struggle to incorporate metadata into their interpretations effectively.
2.2 Data Synthesis and In-Context Learning for Addressing Data Scarcity
Various approaches have been proposed to address the scarcity of table data. Hou et al. (2024) suggested generating synthetic table datasets through AI-driven augmentation techniques. Meanwhile, Wang et al. (2024) introduced Zero-shot Prompting and Few-shot Prompting as efficient alternatives for adapting LLMs to table tasks with minimal data. However, studies have shown that repeated training on synthetic datasets can degrade the quality and diversity of the data, leading to performance deterioration Alemohammad et al. (2024).
Few-shot prompting is a method where examples, explanations, and instructions are provided as input to guide the model’s responses without directly updating its parameters Brown et al. (2020). This technique eliminates the need for retraining the model, making it particularly useful in environments with limited datasets. In this study, we employ this approach to generate interpretations for segments of tables while minimizing computational costs.
@@ 3 Tabular-TX Pipeline
3.1 Theme-Explanation Structure
We propose Tabular-TX, a pipeline for Theme-Explanation structure-based table summarization, designed to generate interpretations for table segments. Unlike conventional approaches that treat table summaries as isolated text generation tasks, our method ensures structural consistency by explicitly organizing content into a Theme Part and an Explanation Part.
3.1.1 Theme Part
The Theme Part serves as a crucial contextual anchor, ensuring that numerical or categorical values in the table are interpreted correctly. It is structured as an adverbial phrase, combining the noun phrase of the table title (table_title) with a citation or basis expression. This structure is essential because the table title provides the sole comprehensive context in table summaries. Unlike general text summaries, table cells alone are insufficient to provide meaningful context, making the resulting sentence ambiguous without additional background information.
3.1.2 Explanation Part
3.2 Sentence Generation Process
@@
Generating sentences with a theme-explanation structure in Tabular-TX involves multiple processing steps to transform tabular data into structured natural language summaries.
3.2.1 Data Transformation
3.2.2 Chain-of-Thought (CoT)
After data transformation, the process moves to generate interpretations for table segments using Chain-of-Thought (CoT) reasoning. This technique is particularly effective for table summarization as it systematically decomposes multi-step reasoning tasks, ensuring that relationships between numerical values and categorical data are captured accurately.
One major issue in table summarization is the Compositional Deficiency problem Zhao et al. (2024), where individual data points are analyzed separately without being adequately integrated into a holistic interpretation. Even if an LLM correctly identifies each table cell individually, the final summary may still lack contextual completeness if the relationships among the values are not recognized.
To mitigate this issue, CoT guides LLMs through a structured reasoning process that progressively refines the interpretation of highlighted table segments. First, the model classifies the type of information contained in the highlighted cells, distinguishing between monetary values, percentages, categorical data, or textual explanations. This step prevents errors, such as misinterpreting percentages as simple numbers, which could yield incorrect results during comparative analysis.
Based on the classification, it selects the most appropriate analytical method, such as enumeration for independent values, magnitude comparison for ranking numerical data, or trend analysis for detecting changes over time. For example, when processing monetary data, the model ensures proper unit conversion to maintain consistency across the summary. Similarly, when dealing with percentage values, a decimal-to-percentage transformation is applied to improve clarity. In cases where the highlighted cells contain free-text explanations, the model leverages background knowledge to provide a more context-aware interpretation.
Afterward, the most suitable analytical method for explaining the relationships between the data is selected and applied to the Explanation Part. For example, enumeration is used if the data items are individually critical, magnitude comparison is used to rank by value, and trend analysis is chosen to track changes over time. These decisions impact the summary’s informativeness.
By integrating these steps logically, CoT ensures that the final summary accurately captures relationships between data points while maintaining structural coherence and readability.
3.2.3 Journalist Persona for Structured Generation
Figure 6: An example of how a persona-applied prompt generates more accurate summaries. (Gray text indicates the English translation.)
@@ 3.2.3 Journalist Persona for Structured Generation
Through strategic prompt engineering, we assign a journalist persona to the LLM to generate Theme-Explanation structured summaries. This persona is particularly effective because table summaries share key characteristics with straight news articles, which prioritize conciseness, objectivity, and fact-based clarity. Rather than generating overly detailed or speculative content, the model produces well-structured and neutral summaries that adhere to journalistic reporting conventions when guided by this persona.
@@
Figure 6 demonstrates the impact of this persona on table summarization. When a generic prompt is used, the model generates an unstructured and somewhat ambiguous summary that captures the core information but lacks contextual clarity and fails to integrate key elements cohesively. Even with specific instructions, the generic prompt still lacks fluency and focuses solely on the facts. In contrast, with the journalist persona applied, the model produces a structured, precise, and contextually enriched summary. This transformation occurs because the journalist persona naturally guides the model to state the source of information explicitly, clearly define numerical constraints, and incorporate additional contextual details that enhance readability. Instead of merely listing numerical values, the model provides a structured sentence that mirrors how financial or administrative data would be reported in news articles.
@@
kobart-base-v2 - Fine-tuned 0.37 0.28 0.35 0.33
@@
Table 1: Evaluation metrics scores for each model.
4 Experimental Setup
4.1 Dataset and Evaluation Metrics
For training and evaluation, we utilized the table interpretation corpus provided by the National Institute of Korean Language NIKL (2024), which focuses on summarizing highlighted table segments into coherent sentences. An example of the dataset is shown in Figure 2. The dataset consists of 7,170 training tables, 876 validation tables, and 876 test tables. Each data point contains metadata such as the document title, table title, publication date, publishing organization, table source URL, highlighted cell information, table data, and a reference summary sentence describing the highlighted portions’ key contents.
@@ 4.1 Dataset and Evaluation Metrics
4.2 Models
4.2.1 Tabular-TX Pipeline Application
To evaluate the effectiveness of the Tabular-TX pipeline, We utilize EXAONE 3.0 7.8B An et al. (2024) and llama-3-Korean-Bllossom-8B [^1] https://huggingface.co/MLP-KTLim/llama-3-Korean-Bllossom-8B models as base models. EXAONE 3.0 7.8B, a successor to EXAONE-LM-v1.0, has demonstrated state-of-the-art performance in Korean TableQA, ranking first on the KorWikiTableQuestions Jun et al. (2022). Similarly, llama-3-Korean-Bllossom-8B is the top-performing sub-10B model in a Korean multi-domain reasoning benchmark. We compare the performance of these models with and without Tabular-TX, assessing whether structured generation (Theme-Explanation format) enhances performance in table summarization. Additionally, we analyze whether Tabular-TX reduces reliance on extensive fine-tuning while maintaining high-quality summaries.
@@ 4.2.2 Additional Adaptation Approaches
In-Context Learning
We begin by applying In-Context Learning (ICL) to each model, providing a few table-summarization examples without explicit fine-tuning. This approach tests how effectively the model can generate coherent sentences for highlighted table cells based solely on a small set of demonstrations.
Low-Rank Adaptation
Next, we assess the computational efficiency and performance of the Tabular-TX pipeline by introducing LoRA (Low-Rank Adaptation). We apply LoRA to the EXAONE 3.0 7.8B model to see if we can maintain high-quality table summaries with fewer resources.
@@
Finally, we evaluate the gogamza/kobart-base-v2 model under a full model fine-tuning setup to determine whether a smaller-scale language model can achieve comparable performance when all its parameters are trained on the table interpretation corpus.
5 Experimental Results
Table 1 presents the performance of various models evaluated using ROUGE-1, ROUGE-L, BLEU, and their average scores. The kobart-base-v2 recorded an average score of 0.33 after fine-tuning. In contrast, EXAONE 3.0 7.8B achieved 0.12 with the ICL method, 0.17 after fine-tuning, and 0.45 when combined with the Tabular-TX method. Similarly, llama-3-Korean-Bllossom-8B, which was also tested with Tabular-TX, showed a notable improvement, reaching an average score of 0.43. These results demonstrate that Tabular-TX consistently outperforms alternative methods, achieving the highest overall performance across different model configurations.
The performance gap between EXAONE 3.0 7.8B and kobart-base-v2, despite both being fine-tuned on the same dataset, can be explained through the multiplicative joint scaling law Zhang et al. (2024a). This principle suggests that when the dataset size is insufficient relative to the model size, the performance gains from fine-tuning remain limited. Since kobart-base-v2 has 124M parameters, while EXAONE 3.0 is approximately 63 times larger, the dataset required to achieve a comparable performance boost must be proportionally scaled up by a factor of 63. The inability to meet this scaling requirement explains why kobart-base-v2’s fine-tuned performance plateaued, while EXAONE 3.0 7.8B demonstrated more significant gains with the same dataset.
@@ 6 Conclusion
This study introduced the Theme-Explanation Table Summarization (Tabular-TX) pipeline, a novel approach to improve table summarization tasks with low-resource requirements.
Experimental results signifies that Tabular-TX enhances table summarization performance. This study contributed to the summarization of complex table data by introducing a novel sentence generation method based on the theme-explanation structure. Furthermore, Tabular-TX achieved excellent performance without fine-tuning, by incorporating ICL. This indicates its potential as a significant contribution to table data analysis, even in resource-constrained environments, without requiring direct model training.
@@
Appendix A Tabular-TX Construction Details

### v3

@@ Abstract
Tables are a primary medium for conveying critical information in administrative domains, yet their complexity hinders utilization by Large Language Models (LLMs). This paper introduces the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline, a novel approach designed to generate highly interpretable summaries from tabular data, with a specific focus on Korean administrative documents. Current table summarization methods often neglect the crucial aspect of human-friendly output. Tabular-TX addresses this by first employing a multi-step reasoning process to ensure deep table comprehension by LLMs, followed by a journalist persona prompting strategy for clear sentence generation. Crucially, it then structures the output into a Theme Part (an adverbial phrase) and an Explanation Part (a predicative clause), significantly enhancing readability. Our approach leverages in-context learning, obviating the need for extensive fine-tuning and associated labeled data or computational resources. Experimental results show that Tabular-TX effectively processes complex table structures and metadata, offering a robust and efficient solution for generating human-centric table summaries, especially in low-resource scenarios.
Theme-Explanation Structure for Table Summarization using
Large Language Models: A Case Study on Korean Tabular Data
TaeYoon Kwack1thanks: Both authors contributed equally to this work.  Jisoo Kim1∗ Ki Yong Jung1 DongGeon Lee2 Heesun Park1thanks: Corresponding author. 1Sungkyunkwan University  2Pohang University of Science and Technology {njj05043, clrdln, wjdrldyd0213}@g.skku.edu donggeonlee@postech.ac.krhspark20@skku.edu
@@ 1 Introduction
Tables are essential for presenting core information, especially within the administrative domain, where critical data is frequently structured in tabular formats Musumeci et al. (2024). The ability of Large Language Models (LLMs) to accurately summarize and elucidate the contents of these tables is becoming increasingly significant for data utilization. An important aspect of effective table summarization is the generation of human-understandable output. This necessitates not only the LLM’s profound comprehension of the input table, but also the crafting of summaries that are both intuitive and concise, delivering key information without ambiguity.
Despite the critical need for human-centric summaries, recent research in table-to-text generation has often overlooked this aspect. Many existing approaches prioritize other metrics or model architectures, without sufficiently addressing how the generated text will be perceived and understood Liu et al. (2024); Zhang et al. (2024b). Consequently, while the output can be factually correct, they lack the clarity, conciseness, or intuitive structure that facilitates effortless human comprehension, particularly when dealing with specialized data such as Korean administrative tables NIKL (2024).
To address this gap and emphasize the generation of human-friendly summaries, we propose the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline. Our approach is meticulously designed to guide LLMs towards producing summaries that are not only accurate but also exceptionally interpretable. First, to ensure a deep understanding of the input table, we decompose the LLM’s reasoning process into a multi-step procedure, where each step focuses on a specific inferential task, thereby simplifying the complex table interpretation process. Second, we employ a “Journalist Persona” prompting strategy to encourage the generation of clear, objective, and well-phrased sentences. Finally, and most critically, we transform these generated insights into a highly structured Theme-Explanation (TX) format, where each summary segment consists of a thematic adverbial phrase followed by a predicative explanatory clause, enhancing readability and directness.
@@
The Tabular-TX pipeline offers significant advantages, particularly in resource-constrained environments where extensive fine-tuning is unfeasible. Our primary contributions are threefold:
  • We introduce a novel pipeline that leverages multi-step reasoning and a journalist persona to generate high-quality textual explanations from complex tabular data.
  • We propose the Theme-Explanation (TX) sentence structure, a new format for table summaries designed to maximize human interpretability.
  • Through empirical evaluation, we demonstrate that our In-Context Learning (ICL)-based approach enables LLMs to achieve strong performance in table data processing and summarization without the need for task-specific fine-tuning.
@@ 2 Related Work
Table-to-Text Generation
@@
TableLlama Zhang et al. (2024b) aims to generalize table-based models beyond task-specific constraints by fine-tuning 14 datasets across 11 tasks, including Highlighted Cells question-answering. The model achieved performance comparable to or surpassing task-specific models and even outperformed GPT-4 OpenAI (2023) on unseen tasks. However, despite its effectiveness, the model suffers from high computational costs.
Among table-based interpretation benchmarks, FeTaQA Nan et al. (2022) serves as a key reference dataset. While Chain-of-Table and TableLlama utilize in-context learning Brown et al. (2020) and fine-tuning, respectively, they struggle to incorporate metadata into their interpretations effectively.
3 Theme-Explanation Structure
Unlike conventional approaches that treat table summaries as isolated text generation tasks, our method ensures structural consistency by explicitly organizing content into a Theme Part and an Explanation Part.
3.1 Theme Part
The Theme Part serves as a crucial contextual anchor, ensuring that numerical or categorical values in the table are interpreted correctly. It is structured as an adverbial phrase, combining the noun phrase of the table title (table_title) with a citation or basis expression [^1] Comparison of summarization sentence with and without theme part is illustrated at Figure 4.. This structure is essential because the table title provides the sole comprehensive context in table summaries. Unlike general text summaries, table cells alone are insufficient to provide meaningful context, making the resulting sentence ambiguous without additional background information.
3.2 Explanation Part
@@ 4 Tabular-TX Pipeline
Generating summaries with the theme-explanation structure in Tabular-TX involves multiple processing steps to transform tabular data into structured natural language summaries. [^2] Within Tabular-TX, data is preprocessed before using LLMs. Further details on data preprocessing are discussed in Appendix C.1.
4.1 Chain-of-Thought (CoT)
After data transformation, the actual sentence generation process begins using Chain-of-Thought (CoT) reasoning. Large Language Models (LLMs) generally face performance degradation when handling tabular data summarization, as this task simultaneously requires multiple capabilities such as table recognition, mathematical reasoning, and commonsense inference. This issue, known as the Compositional Deficiency problem Zhao et al. (2024), occurs because individual data points tend to be analyzed separately without adequately integrating their relationships into a holistic interpretation. CoT mitigates this by systematically guiding the model to tackle one reasoning step at a time, thereby improving interpretative accuracy and contextual completeness.
Specifically, CoT first classifies the types of highlighted cells, distinguishing among monetary values, percentages, categorical data, and textual explanations. This classification step prevents potential errors, such as misinterpreting percentages as plain numbers. Then, depending on the classified data type, the most appropriate analytical method—such as enumeration for listing individual items, magnitude comparison for numerical rankings, or trend analysis for temporal changes—is selected and applied. For instance, monetary values are converted into consistent units, and percentages are appropriately formatted for clarity.
In the context of Korean administrative table data, these challenges are further complicated by the language’s implicit nature, the potential gap between administrative and everyday terminology, and morphological complexities (e.g., absent subjects or ambiguous particle usage). CoT systematically decomposes these linguistic hurdles by (1) classifying specialized terms, (2) normalizing numeric expressions in line with Korean usage conventions, and (3) incrementally integrating contextual cues, such as clarifying administrative vocabulary or disambiguating omitted referents. Through this stepwise process, the model avoids misinterpretations caused by either unfamiliar terms or implicit structures, ultimately generating summaries that better align with Korean textual norms.
By addressing each reasoning subtask explicitly and sequentially, CoT ensures the final table summary captures data relationships clearly and coherently, resulting in an accurate and contextually meaningful summary.
4.2 Journalist Persona for Structured Generation
@@
In last step of pipeline we assign a journalist persona to the LLM to generate Theme-Explanation structured summaries. This persona is particularly effective because table summaries share key characteristics with straight news articles, which prioritize conciseness, objectivity, and fact-based clarity. Rather than generating overly detailed or speculative content, the model produces well-structured and neutral summaries that adhere to journalistic reporting conventions when guided by this persona.
@@
Figure 2 demonstrates the impact of the journalist persona on table summarization. With a generic prompt, the model generates an ambiguous summary that captures core information but lacks contextual clarity and coherence. In contrast, applying the journalist persona produces a structured and contextually enriched summary. This improvement occurs because the journalist persona explicitly guides the model to state information sources, clearly define numerical constraints, and incorporate contextual details, closely resembling news article formats.
@@
KoBART - Fine-tuned 0.37 0.28 0.35 0.33
@@
Table 1: Evaluation results on the Korean Korean table interpretation benchmark for each model.
5 Experiment
5.1 Experimental Setup
Dataset
For training and evaluation, we utilized the Korean table interpretation benchmark NIKL (2024), which focuses on summarizing highlighted table segments into coherent sentences. An example of the dataset is shown in Figure 3. The dataset consists of 7,170 training tables, 876 validation tables, and 876 test tables. Each data point contains metadata such as the document title, table title, publication date, publishing organization, table source URL, highlighted cell information, table data, and a reference summary sentence describing the highlighted portions’ key contents.
Evaluation Metrics
@@ 5.1 Experimental Setup
Models
To evaluate the effectiveness of the Tabular-TX pipeline, we utilize EXAONE 3.0 7.8B An et al. (2024) and llama-3-Korean-Bllossom-8B [^3] https://huggingface.co/MLP-KTLim/llama-3-Korean-Bllossom-8B models as base models. EXAONE 3.0 7.8B, a successor to EXAONE-LM-v1.0, has demonstrated state-of-the-art performance in Korean TableQA, ranking first on the KorWikiTableQuestions Jun et al. (2022). Similarly, llama-3-Korean-Bllossom-8B is the top-performing sub-10B model in a Korean multi-domain reasoning benchmark. We compare the performance of these models with and without Tabular-TX, assessing whether structured generation enhances performance in table summarization. Additionally, we analyze whether Tabular-TX reduces reliance on extensive fine-tuning while maintaining high-quality summaries.
@@ 5.1.1 Additional Adaptation Approaches
In-Context Learning (ICL)
We begin by applying ICL to each model, providing a few table-summarization examples without explicit fine-tuning. This approach tests how effectively the model can generate coherent sentences for highlighted table cells based solely on a small set of demonstrations.
Low-Rank Adaptation (LoRA)
Next, we assess the computational efficiency and performance of the Tabular-TX pipeline by introducing LoRA Hu et al. (2022). We trained the EXAONE 3.0 7.8B by applying LoRA to see if we could maintain high-quality table summaries with fewer resources.
@@
Finally, we evaluate the KoBART (Korean BART) [^4] https://huggingface.co/gogamza/kobart-base-v2 model under a full model fine-tuning setup to determine whether a smaller-scale language model can achieve comparable performance when all its parameters are trained on the Korean table interpretation benchmark.
5.2 Experimental Results
Table 1 presents the performance of various models evaluated using ROUGE-1, ROUGE-L, BLEU, and their average scores. The KoBART recorded an average score of 0.33 after fine-tuning. In contrast, EXAONE 3.0 7.8B achieved 0.12 with the ICL method, 0.17 after fine-tuning, and 0.45 when combined with the Tabular-TX method. Similarly, llama-3-Korean-Bllossom-8B, which was also tested with Tabular-TX, showed a notable improvement, reaching an average score of 0.43. These results demonstrate that Tabular-TX consistently outperforms alternative methods, achieving the highest overall performance across different model configurations.
The performance gap between EXAONE 3.0 7.8B and KoBART, despite both being fine-tuned on the same dataset, can be explained through the multiplicative joint scaling law Zhang et al. (2024a). This principle suggests that when the dataset size is insufficient relative to the model size, the performance gains from fine-tuning remain limited. Since KoBART has 124M parameters, while EXAONE 3.0 is approximately 63 times larger, the dataset required to achieve a comparable performance boost must be proportionally scaled up by a factor of 63. The inability to meet this scaling requirement explains why the performance of the fine-tuned KoBART plateaued, while EXAONE 3.0 7.8B demonstrated more significant gains with the same dataset.
@@ 6 Conclusion
This study introduced the Theme-Explanation Table Summarization (Tabular-TX) pipeline, a novel approach to improve table summarization tasks with low-resource requirements. Moreover, the proposed pipeline effectively overcame unique challenges in Korean administrative table processing.
Experimental results signify that Tabular-TX enhances table summarization performance. This study contributed to the summarization of complex table data by introducing a novel sentence generation method based on the theme-explanation structure. Furthermore, Tabular-TX achieved excellent performance without fine-tuning, by incorporating ICL. This indicates its potential as a significant contribution to table data analysis, even in resource-constrained environments, without requiring direct model training.
@@
Appendix A Korean table interpretation benchmark
We leverage the Korean table interpretation benchmark provided by the National Institute of Korean Language NIKL (2024).
As shown in Figure 3, an objective of the dataset is to summarize the highlighted cells, which are labeled in the metadata as highlighted_cells, into a single coherent sentence.
Appendix B Examples of the Theme Part & Explanation Part
B.1 Theme Part
B.2 Explanation Part
Appendix C Implementation Details
C.1 Data Preprocessing
C.2 Prompt Details